OpenStudy (keigh2015):

The length of a rectangle is represented by the function L(x) = 6x. The width of that same rectangle is represented by the function W(x) = 4x2 − 3x + 8. Which of the following shows the area of the rectangle in terms of x?

OpenStudy (keigh2015):

@misssunshinexxoxo

OpenStudy (keigh2015):

(L ⋅ W)(x) = 24x3 − 18x2 + 48x (L ⋅ W)(x) = 24x3 − 3x + 8 (L + W)(x) = 4x2 + 3x + 8 (L + W)(x) = 4x2 − 9x + 8

OpenStudy (misssunshinexxoxo):

Which one do you believe it is?

OpenStudy (keigh2015):

I think the first one

OpenStudy (misssunshinexxoxo):

We know the length is 6 "W(x) = 4x2 − 3x + 8" the first is correct after plugging it in

OpenStudy (misssunshinexxoxo):

All you got to do with area is plug it in and sometimes factor; this is leading you into geometry. Get ready for it.

OpenStudy (keigh2015):

So I multiply? This is getting me to confused

OpenStudy (misssunshinexxoxo):

Yes, length x width

OpenStudy (keigh2015):

ok so did i have it correct? with the first one?

OpenStudy (misssunshinexxoxo):

Solve this for the width W(x) = 4x2 − 3x + 8; Yes you did :)

OpenStudy (keigh2015):

awesome thank you so much

OpenStudy (haseeb96):

A is the correct answer btw guys and you did well :) my boy @Keigh2015